b.     Given the population 2,1,3,3,5,4,7 find the probability that a random sample of size 25, selected with replacement, will yield a sample mean greater than 3.7, but less than 4.3.



1
Expert's answer
2021-05-24T19:00:12-0400

Given population is- (2,1,3,3,5,7,4)


population mean "\mu=\dfrac{2+1+3+3+5+4+7}{7}=\dfrac{25}{7}=3.55"


Sample standard deviation-


"s=\sqrt{\dfrac{(x-\mu)^2}{n}}"


"=\sqrt{\dfrac{(-1.55)^2+(-2.55)^2+(-0.55)^2+(-0.55)^2+(1.45)^2+(0.45)^2+(3.45)^2}{7-1}}"


"=\sqrt{\dfrac{23.7175}{7}}=\sqrt{3.3882}=1.9892"


Required Probability


"P(3.7<x<4.3)=P(\dfrac{3.7-3.77}{1.9892}<z<\dfrac{4.3-3.77}{1.9892})"


"=P(-0.035<z<0.266)\\[9pt]=0.11884"